# Multi-view Structural Convolution Network for Domain-Invariant Point Cloud Recognition of Autonomous Vehicles

### v3

@@ 5.5 Ablation study
To evaluate the impact of our unseen domain generalization methods on improving the leaning of domain-invariant feature representation, we performed various experiments based on both intra-domain and cross-domain classification, which are illustrated in Table 8. MSCN, which can increase the amount of training without overfitting, performs better than MSCN for single-domain classification. Note that MSCN does not include a domain generation module. With the domain generation module, MSCN demonstrates even more robust performance than MSCN in various domain change scenarios, with an average improvement of 5.8%, indicating that our proposed unseen domain generation method ensures domain generalization ability.
Figure 6: Evaluation of Invariance Properties. All models were trained on the PanKyo [30] dataset and then tested on (a) PanKyo[30], (b) KITTI[25], and (c) nuScenes[26] datasets. The graphs on the left illustrate the performance of the models when point clouds were rotated along the upward direction. The graphs on the right show the results when the models were tested with randomly shifted point clouds within a certain distance in all directions. For example, the right graph in (a) presents the results of 11 experiments (from 0 to 100 shift) conducted in an Intra-domain setting. Each experiment measures the performance of the models when the point clouds are randomly shifted within a specified maximum distance range.
Finally, we compare the robustness of our model in domain shift scenarios, as well as its invariance properties when point clouds undergo rotations and shifts, against other models. To evaluate this, we progressively rotated the point clouds from 0 (no rotation) to 100 in 10 increments and randomly shifted the point clouds from 0 to 100 units.
Figure 6 illustrates six different scenarios where point clouds were subjected to both shifts and rotations in intra-domain and cross-domain settings. In the intra-domain experiments, as the point clouds were gradually rotated, most of the methods showed a sharp decrease in performance, while our MSCN and MSCN exhibited a more gradual decrease, demonstrating their resilience. Furthermore, when point clouds were shifted within the intra-domain setting, our MSCN and MSCN again proved to be the most robust.
Furthermore, in domain shift scenarios, including PanKyo to KITTI and PanKyo to nuScenes, our models outperformed others, maintaining strong performance even as the point clouds were rotated and shifted. This shows that MSCN and MSCN are not only effective in maintaining accuracy, but are also robust to significant geometric transformations in challenging cross-domain environments.
6 CONCLUSION
@@
In this paper, we introduced MSCN, which learns domain-invariant feature representations by capturing the geometric information of 3D point clouds from the standpoint of both local context and overall context. Moreover, we developed the MSCN through our unseen domain expansion framework, enabling the MSCN to learn more robust feature representations. Our MSCN and MSCN demonstrated the highest performance in cross-domain classification tasks in self-driving data sets, which are derived not only from real-world data, different regions, and sensors, but also from simulators. These results indicate that our method addresses sensor-dependent problems in point cloud classification tasks within the self-driving vehicle field. In light of these findings, our future research endeavors will focus on expanding our method to object detection tasks for real-world applications that require domain-invariant recognition, such as autonomous driving, robotics, and virtual reality.
@@ Acknowledgment
This work was supported by the Korea National Police Agency(KNPA) under the project ’Development of the autonomous driving patrol service for active prevention and response to traffic accidents’ (RS-2024-00403630).

### v4

@@ V-E Ablation study
To evaluate the impact of our unseen domain generalization methods on improving the leaning of domain-invariant feature representation, we perform various experiments based on both intra-domain and cross-domain classification, which are illustrated in Table VIII. MSCN, which can increase the amount of training without overfitting, performs better than MSCN for single-domain classification. Note that MSCN does not include a domain generation module. With the domain generation module, MSCN demonstrates even more robust performance than MSCN in various domain change scenarios, with an average improvement of 5.8%, indicating that our proposed unseen domain generation method ensures domain generalization ability.
VI CONCLUSION
@@
In this paper, we introduced MSCN, which learns domain-invariant feature representations by capturing the geometric information of 3D point clouds from the standpoint of both local context and overall context. Moreover, we developed the MSCN through our unseen domain expansion framework, enabling the MSCN to learn more robust feature representations. Our MSCN and MSCN demonstrated the highest performance in cross-domain classification tasks across all AV datasets, which are derived not only from real-world data, different regions, and sensors, but also from simulators. These results indicate that our method addresses sensor-dependent problems in point cloud classification tasks for the deployment of AVs. In light of these findings, our future research endeavors will focus on expanding our method to object detection tasks for AV systems that require domain-invariant recognition.
@@ Acknowledgment
This work was supported by the Korea National Police Agency(KNPA) under the project ‘Development of the autonomous driving patrol service for active prevention and response to traffic accidents’ (RS-2024-00403630).